# Uncertainty Quantification of Surrogate Models using Conformal Prediction

### v2

@@ 3.3.1 U-Net
Figure 7: Calibrated prediction sets obtained by performing conformal prediction with the three nonconformity scores mentioned in Section 2.2 for the 2D wave equation. The color bar depicts the coverage (in %) provided by each prediction set, characterised by the $1-\alpha$ value. Each figure shows a slice of the prediction set and exact inference value taken along the $y$-axis for the final time instance of the prediction. The exact numerical solution (black with white markers) can be seen to fit up to the coverage offered by the prediction set.
U-Net is designed to perform a single feed-forward map, taking in 20-time instances of the field to produce an output of shape $[30,33,33]$, characterised by 30-time instances across a $33\times 3$ 33 spatial domain. A calibration and validation dataset comprising 100 simulations each is generated from the same distribution as that of the training data to evaluate CP for the wave equation. Coverage is estimated at each point in the output tensor as given in Section 2.1.2.
Comparing the different methods used for CP (in Figure 3 and Table 1), we see that the AER method is the least computationally intensive and offers valid coverage. This is further accentuated by Figure 7, where the tightness of the fit and the coverage obtained by performing CP at various values of $\alpha$ is given. All methods provide valid coverage in tandem with the ideal coverage as expected by Equation 1, where those using quantile regression and AER errors provide overly conservative error bars. This is expected as Equation 1 offers an inequality, offering wide predictive sets for the specific value.
In Figure 7, we take the slice along the $y$-axis of the prediction sets at various $\alpha$ levels to visualise the uncertainty of the predicted field for that slice. The slices represent the variation of the field along the y-axis for the final time instance of the prediction. Figure 7 shows that the tighter fits are obtained using the AER and STD methods, where the quantile regression fails to provide a tight fit although it provides good coverage.
To further iterate the impact of the CP framework as a method of uncertainty quantification, we further generate calibration and validation data, but this time from a different physical model. The newer datasets are sampled by solving the wave equation with wave velocity at half the speed of that used for the training data. This allows us to further benchmark the CP formulation in the paper in cases where the calibration and prediction regimes are outside of the training distribution. As given in Figure 2, when utilised outside of the training domain, the uncertainty captured by dropout (STD) fails to capture the modelling failure. However, by calibrating the uncertainty using the CP framework, it can capture the model’s (in)capabilities with statistical guarantees, irrespective of the training conditions of the pre-trained model. This becomes particularly important when we consider the utility of the surrogate model in out-of-training-distribution scenarios, where we will need more data to fine-tune the network.
@@ 3.3.2 FNO
Furthering the experiment’s complexity, FNOs are trained in an autoregressive framework to model the spatio-temporal evolution of the wave. The autoregressive model takes in the first 20 time instances to output the next 10 time instances, further unrolled for the next 60 time instances. We perform conformal prediction across the entire time rolled-out output of shape $[60,33,33]$. Since the FNO offers the best performance when trained using a relative LP norm, we did not perform CQR for experiments using the FNO.
While performing conformal prediction over the FNO, we also demonstrate its utility in obtaining statistically guaranteed coverage for inference over a physics scenario different from the one on which the FNO was trained on as mentioned in the earlier section (also see Appendix C). The FNO is applied in the same in-distribution and out-of-distribution setting as the U-net. Through this, we demonstrate the validity of the error bars even when predicting out-of-training-distribution data.
Figure 23 (in the appendix) shows that we obtain guaranteed coverage for predictions across calibration methods, even when tested outside the training distribution. As shown in Figure 8, the CP framework provides valid error bars for the desired coverage, irrespective of the training conditions. The only criterion that needs to be met is that the prediction regime and the calibration regime are exchangeable. The inductive CP framework can provide valuable uncertainty quantification, even when deployed to solve for those unseen sets of solutions. Furthermore, FNO offers a tighter fitting coverage than the U-Net used for the wave equation; see Figure 3.
Figure 8: Visualising the errors obtained using the CP framework for an auto-regressive FNO within (Figure (a)) and outside the training distribution (Figure (b)). In both figures, the top row represents the numerical solution of the wave equation, the FNO prediction in the middle row and the calibrated error for 90 % coverage.
@@
The 2D Navier–Stokes equations are a family of PDEs that describe the motion of viscous fluids. The equations establish a mathematical model for the conservation of mass, momentum, and energy in a fluid flow. Considering its complexity and strong non-linearity the Navier–Stokes equations are solved using computational fluid dynamics (CFD). Neural-PDEs have emerged as a suitable method to assist complex CFD simulators to solve the Navier–Stokes equations at scale (Azizzadenesheli et al., 2024). Within this experiment, an FNO is trained to model the evolution of vorticity as described in the original FNO paper (Li et al., 2021). Further descriptions of the physics formulation, datasets and training can be found in Appendix D. The model was trained from simulations with viscosity $\nu=1e-3$ and then calibrated to form the prediction set for viscosity $\nu=1e-4$. The datasets were gathered from the original FNO paper Li et al. (2021). The coverage obtained across the out-of-distribution prediction sets can be seen in Figure 3. We modify the FNO in (Li et al., 2021) with dropout layers, to train a probabilistic neural operator. As shown in Figure 9, deploying the CP framework across the probabilistic outputs to be further calibrated to obtain guaranteed coverage.
Figure 9: Calibrating the uncertainty captured by using MC dropout (STD) within the FNO in modelling out-of-distribution data for the Navier–Stokes case. The top row shows the ground truth, the second row the output of the FNO, the third row the error (taken as the standard deviation here) captured by the probabilistic FNO, and the final row shows the calibrated error obtained using the CP framework over the probabilistic outputs showing 67 % coverage.
@@
Magnetohydrodynamics (MHD) governs the evolution of a plasma state within a fusion device, such as a Tokamak. It is characterised by the coupling of Navier–Stokes equation of fluid dynamics and Maxwell’s equations of electromagnetism (Bellan, 2006). The evolution of MHD fields involves modelling multi-physics systems with coupled variables $(\theta)$ that are integrated together. We model the evolution of multiple plasma blobs in a non-uniform temperature field with a reduced-MHD model as described by (Hoelzl et al., 2021). The model governs the spatio-temporal evolution of the density ($\rho$), electrostatic potential ($\Phi$) and temperature ($T$) within a simple circular geometry, modelled in the toroidal coordinates $R$ and $Z$. The full description of the MHD physics we are interested in modelling is provided in Appendix E.
Figure 10: Spatial slice plots showing 90 percent coverage ($\alpha=0.1$) over the multi-variable FNO predictions, calibrated using conformal prediction. The figures represent the spatial distribution of the plasma density $\rho$ (left column), electrostatic potential $\Phi$ (middle column) and plasma temperature $T$ (right column) along the axes as obtained from the numerical solution, FNO and the upper and lower bands at the $20^{th}$ timestep. The top plots portray the field distribution along the $R$-axis and the bottom plots the field distribution along the $Z$-axis. Across the dimensions of the output, we can obtain valid error bars, fitting within the desired coverage for our surrogate model.
The dataset and the trained model were taken from (Gopakumar et al., 2024). The data was constructed by running simulations of the reduced-MHD cases using the code JOREK (Hoelzl et al., 2021). The physics of the equation was kept stationary throughout the campaign; however, the initial conditions that characterise the blobs were changed within each simulation (see Appendix E). A total of 2000 simulations/data points were generated, out of which 1000 were used for training, 500 for calibration and 500 for validating the prediction sets.
[…]
@@ 3.6 Foundational Physics Models
Foundation models (Bommasani et al., 2022) capable of performing multi-task modelling across a range of PDE-related tasks have been developed in the recent past (McCabe et al., 2023; Alkin et al., 2024; Hao et al., 2024; Rahman et al., 2024). These models rely on deploying novel transformer-based architectures that deploy an attention mechanism across the spatio-temporal domain to pre-train across various families of PDEs, allowing complex multi-physics scenarios to be modelled. The underlying principle relies on the idea that PDEs describing various physical systems, share similar differential operators. The models have shared embeddings across the variables to characterise the different physical operators, such as diffusion or convection. This allows them to learn the global behaviour governing different PDEs, leaving the local features to be learnt in the fine-tuning stage Alkin et al. (2024).
As these models scale and find further application across a range of domains, it becomes increasingly critical to quantify the uncertainty associated with the model predictions. Being pre-trained on diverse PDE datasets, these models often find applications in similar tasks. Considering the complexity of scale and the importance of the task, it is hard to validate the utility of these pre-trained models in an application without being able to provide UQ estimates that can be guaranteed. The same extends to fine-tuned models, trained for a specific downstream application. By performing conformal prediction across the model outputs at each space, we are able to validate the model performance for a specific application. CP relies on requiring ground truth data to help provide statistical validity, enforcing some computational effort in generating the numerical data. However, in the case of the application of a fine-tuned model for a specific application, calibration data already exists as the fine-tuning data. The central utility of this CP framework within the domain of surrogate modelling becomes clearer when extended to the fine-tuning setting where the prediction domain is the same as the training domain, reducing the need for further collection of data for calibration.
Within this experiment, we restrict our focus on applying the CP framework across one of the early attempts at a Foundational Physics Model, demonstrated in (McCabe et al., 2023). The authors modify the architecture of an adaptive vision transformer (AViT) (Yin et al., 2022a) with shared embedding and normalisation strategy across the variable space. The model uses an AViT backbone which attends over the space and time axis sequentially. The model is trained autoregressively to output the next time instance of the field(s) across the spatial domain. For further details on the model architecture, training setup, and the datasets used for pre-training, we refer to the original paper (McCabe et al., 2023).
@@ 3.6.1 Pre-trained Model: Zero-Shot Learning
Within the first experiment, the efficacy of the pre-trained foundational physics model to evolve the unseen physics is explored in conjunction with the formulated CP framework. The largest Multi-Physics Pre-trained (MPP-AVIT-L) from (McCabe et al., 2023), consisting of 409 million parameters, is tested on the MHD dataset as detailed in Section 3.5. The MPP-AVIT-L model is pre-trained to model across a range of equations: shallow-water, diffusion-reaction and Navier–Stokes equations, where the model learnt to model densities, velocities and pressure formulations. Thus, we isolate the density fields from the MHD data to perform inference over the MPP-AVIT-L model. The model is structured to take in the first 16 time steps for each simulation and then output the next time step. This procedure is autoregressively continued until the $50^{th}$ time instance.
Figure 11: Performance of the MPP-AViT model in modelling MHD density (out-of-training-distribution data). The top row shows the numerical solution, the second row shows the model output, the third row shows the absolute error, and the bottom row shows the calibrated error (width of the error bars) providing 95% coverage.
[…]
@@ 3.6.2 Fine-tuned Model
The next iteration involves fine-tuning the MPP-AViT model on the density fields from the MHD dataset to create a foundation model well-versed in magnetohydrodynamics found within the core of a Fusion reactor. We fine-tune the tiny version of the foundational physics model (MPP-AViT-Ti) using $3/4$ of the available data for training and calibration. The remaining $1/4$ is used for evaluation and obtaining the prediction sets. Within the fine-tuning context, the training data doubles up as calibration data. The intention of fine-tuning is often to bring a certain data regime within the distribution of the model, hence the fine-tuned model’s performance is only of interest within the new domain. This allows us to care about performance in the same distribution as that of the fine-tuning data, allowing us to use the same data to calibrate the model as we are only interested in performance within the same domain. Since the objective of fine-tuning is to align a model with a specific task, the usage of training data as calibration data is well justified. From Figure 12, we notice that fine-tuning the foundation model on the MHD model resulted in a significant improvement in performance as compared to the previous instance of zero-shot learning.
Figure 12: Demonstrating the Performance of the Fine-tuned MPP model in capturing the MHD behaviour. The top row demonstrates the numerical solution, the second row shows the model output, the third row shows the absolute error and the bottom row demonstrates the calibrated error providing 95% coverage. Being a deterministic model, the calibration procedure results in generating general uncertainties for regions of major dynamics within the dataset, not specific to the prediction instance.
Though the CP framework offers a coverage guarantee (see Appendix F), the calibrated error values end up being more general, characterising the global dynamics and uncertainties. This arises due to the deterministic nature of the model, not providing input dependence to the prediction sets. For a considerably well-fit, as seen in Figure 12, only the global uncertainties are captured.
@@ 3.7 Neural Weather Prediction
In addition to surrogate models of systems described by explicit PDEs, the proposed methodology is also applicable to more general machine learning models describing physical processes. To demonstrate this, we here study the use of CP for data-driven weather forecasting models. Traditional weather forecasting models typically combine PDEs describing large-scale interactions and parametrizations describing subgrid-scale physical processes (Kalnay, 2002). Data-driven machine learning models approximate this whole process with a single neural network model. This allows for orders of magnitude faster forecasting speed and, when training on data incorporating observations, also more accurate forecasts (Bouallègue et al., 2024; Bi et al., 2023; Kurth et al., 2023; Lam et al., 2023).
Due to the chaotic nature of the weather system, capturing uncertainty in weather forecasts has long been an important consideration both in research and operations. Such probabilistic modeling has typically been achieved by ensemble forecasting, where perturbations are used to produce samples of possible forecast trajectories (Coiffier, 2011). Existing data-driven models are still largely deterministic (Rasp et al., 2024). There are attempts to produce ensemble forecasts using machine learning models by perturbing initial states (Kurth et al., 2023; Chen et al., 2023b), training multiple models (Graubner et al., 2022) or generative modeling (Hu et al., 2023; Price et al., 2023; Oskarsson et al., 2024). Fundamentally, ensemble forecasting always requires a computational cost proportional to the number of ensemble members i.e. the number of forecasts made via perturbations. In contrast, CP offers a cheap method to immediately quantify forecast uncertainty for a time, position, and variable of interest. This uncertainty can be used by meteorologists interpreting the forecast, conveyed to decision-makers reacting to extreme weather events or directly presented to end-user looking up the forecast for the coming week. As CP enables UQ for one single forecast output by the model, it is directly applicable to existing deterministic machine learning models. A limitation of scalar uncertainty estimates is that there are no samples of the distribution over the atmospheric state. In some scenarios it can be valuable to inspect such samples to gain an understanding of how different weather scenarios are unfolding.

### v3

@@ 3.3.1 U-Net
The U-Net performs a single feed-forward mapping from 20 input time steps to 30 output time steps, producing outputs of shape $[30,33,33]$. Coverage is estimated cell-wise across the output tensor following Section 2.1.1.
Figure 7: Prediction sets for the 2D wave equation using three nonconformity scores. Colour bars show coverage levels (in %, corresponding to $1-\alpha$). Each panel displays a spatial slice along the $y$-axis at the final time step. The ground truth (black with white markers) lies within the calibrated bounds, validating Equation 1.
Figure 7 shows spatial slices of prediction sets at multiple $\alpha$ levels. All three methods achieve valid coverage (Figure 3, Table 1), with AER being computationally cheapest. Both CQR and AER produce conservative (wide) intervals, as expected from the inequality in Equation 1. AER and STD provide tighter fits than CQR while maintaining coverage guarantees.
Out-of-Distribution Testing.
To test robustness, we generate new calibration and validation datasets by solving the wave equation with half the wave velocity used during training. This tests CP’s effectiveness when the calibration regime differs from the training distribution. As shown in Figure 2, uncalibrated MC dropout fails to capture modelling errors in this out-of-distribution regime. In contrast, CP provides statistically guaranteed bounds regardless of the model’s training conditions. This demonstrates CP’s value for deploying pre-trained surrogates in new physical regimes without costly retraining.
@@ 3.3.2 FNO
We train an FNO in an autoregressive framework: the model takes 20 initial time steps, predicts the next 10 steps, then recursively unrolls to produce 60 total output steps (shape $[60,33,33]$). CP is performed over the entire rolled-out output. Since FNOs perform best with relative $L^{p}$ loss, we omit CQR for this architecture.
The FNO is tested in both in-distribution and out-of-distribution (half-speed) settings, matching the U-Net experiments. Figure 8 shows that CP provides valid error bars under both conditions. The key requirement is exchangeability between calibration and prediction regimes, not similarity to training data. This enables valuable UQ even for previously unseen solution families. The FNO achieves tighter coverage than the U-Net (Figure 3), likely due to its operator learning formulation. Additional coverage plots are in Figure 23 (Appendix C).
Figure 8: CP performance for autoregressive FNO on the wave equation. Top row: ground truth; middle row: FNO predictions; bottom row: calibrated error bars (90% coverage). Valid coverage is maintained both within (a) and outside (b) the training distribution, demonstrating CP’s robustness to distribution shift.
@@
The incompressible 2D Navier–Stokes equations describe viscous fluid dynamics, modelling conservation of mass and momentum. Their complexity and strong nonlinearity necessitate computational fluid dynamics (CFD) solvers. Neural-PDE methods offer efficient alternatives at scale (Azizzadenesheli2024). Following li2021fourier, we train an FNO to model vorticity evolution. The model is trained on simulations with viscosity $\nu=10^{-3}$, then calibrated and tested on data with $\nu=10^{-4}$ (out-of-distribution). The FNO maps 10 input time steps to the next 10 output steps. Physics details and training specifications are in Appendix D.
To enable STD-based CP, we modify the FNO architecture by adding dropout layers, creating a probabilistic operator. The resulting model outputs both mean predictions and uncertainty estimates via MC dropout sampling.
Figure 9: Calibrating probabilistic FNO predictions for out-of-distribution Navier–Stokes data. Rows show: ground truth, FNO predictions, uncalibrated MC dropout uncertainty (standard deviation), and calibrated CP error bars (67% coverage). The uncalibrated dropout underestimates errors, but CP corrects this to achieve the target coverage.
Figure 9 demonstrates CP’s ability to calibrate probabilistic models. The uncalibrated dropout-based uncertainty provides only 7.52% coverage, severely underestimating prediction errors. CP adjusts these intervals to achieve the target 67% coverage (we show 67% instead of the standard 90% to better visualise the calibration effect at moderate coverage levels). Coverage validation across all $\alpha$ levels is shown in Figure 3. This helps illustrate CP’s dual utility: providing guarantees for deterministic models (via AER) and calibrating uncalibrated probabilistic models (via STD).
@@
Magnetohydrodynamics (MHD) couples the Navier–Stokes equations with Maxwell’s equations to model plasma evolution in fusion devices such as tokamaks (bellan2006fundamentals). We consider a reduced-MHD system (Hoelzl2021jorek) describing multiple plasma blobs in a non-uniform temperature field. The system evolves three coupled variables—density ($\rho$), electrostatic potential ($\Phi$), and temperature (TT)—on a $106\times 106$ toroidal grid (coordinates $R$, $Z$). This represents a challenging multi-variable, multi-physics problem. Full physics equations are in Appendix E.
We use 2000 simulations from the JOREK code (Hoelzl2021jorek), split into 1000 for training, 500 for calibration, and 500 for validation. Each simulation varies the initial conditions (blob positions, widths, amplitudes) while keeping the physics parameters fixed. The dataset and pre-trained model are taken from Gopakumar_2024.
A multi-variable FNO learns the coupled dynamics of all three fields simultaneously. The model autoregressively predicts from 10 input time steps to 5 output steps, recursively continuing to the 50th time step. CP is performed over the complete spatio-temporal domain, yielding prediction sets for each cell in the 4D output tensor (time $\times$ $R$ $\times$ $Z$ $\times$ variables). Given the model’s scale (9.4M parameters, 1.35M output dimensions), we use only the AER method for computational efficiency.
Figure 10: Spatial profiles at timestep 20 showing 90% coverage ($\alpha=0.1$) for multi-variable MHD predictions. Columns: density $\rho$ (left), potential $\Phi$ (center), temperature $T$ (right). Rows show profiles along the $R$-axis (top) and $Z$-axis (bottom). For each variable, we display the numerical solution, FNO prediction, and calibrated upper/lower bounds. CP accurately captures the sharp features of plasma blobs across all three coupled variables.
Figure 10 shows spatial slices through the prediction sets for all three variables at a single time step. The calibrated error bars successfully bound the sharp density, potential, and temperature peaks characterising the plasma blobs. Coverage validation (Figure 3) confirms guaranteed marginal coverage across all 1,348,320 output dimensions ($50\times 106\times 106\times 3$), demonstrating CP’s scalability and immunity to the curse of dimensionality. The bounds provide interpretable confidence estimates for each variable’s spatial distribution, enabling assessment of the surrogate’s reliability across the operational domain.
@@ 3.6 Foundational Physics Models
Foundation models pre-trained on diverse PDE datasets (bommasani2022opportunities; mccabe2023multiple; alkin2024upt; hao2024dpot; rahman2024pretraining) have emerged as a promising approach for multi-task scientific modelling. These models employ transformer-based architectures with attention mechanisms across spatio-temporal domains, learning shared representations of differential operators (e.g., diffusion, convection) common across PDE families. This enables them to capture global behaviours during pre-training, leaving task-specific local features for fine-tuning (alkin2024upt).
As these models scale and deploy across safety-critical applications, UQ becomes essential. CP offers an efficient validation framework: for fine-tuned models, the existing fine-tuning data can serve as calibration data, eliminating the need for additional simulations. This is justified because fine-tuning aims to align the model with a specific target distribution, making performance within that distribution the primary concern.
We apply CP to the Multi-Physics Pre-trained Adaptive Vision Transformer (MPP-AViT) of mccabe2023multiple. This model uses shared embeddings and normalisation across variables, with an AViT backbone (yin2022avit) that sequentially attends over space and time. The model autoregressively predicts the next time step given current field values. We refer to mccabe2023multiple for full architectural and training details.
@@ 3.6.1 Pre-trained Model: Zero-Shot Learning
We test the largest pre-trained model (MPP-AViT-L, 409M parameters) on MHD density evolution—a physics regime not seen during training. The model was pre-trained on shallow-water, diffusion-reaction, and Navier–Stokes equations, learning to model densities, velocities, and pressures. We extract only density fields from our MHD dataset (Section 3.5) for inference. The model takes 16 input time steps and autoregressively predicts a single step forward until the 50th time step.
Figure 11: Zero-shot MPP-AViT performance on out-of-distribution MHD density evolution. Rows show: ground truth, model predictions, absolute error, and calibrated error bar width (95% coverage). Despite no training on MHD physics, the model captures major blob features. CP quantifies the prediction uncertainty with guaranteed coverage.
Figure 11 shows that despite zero exposure to MHD during training, the model captures the major features of plasma blob evolution (radial outward motion). However, finer details are lost, and patch-based artefacts appear, a known limitation of the architecture (mccabe2023multiple). Using 1000 calibration data points and the AER method, CP provides valid 95% coverage bounds (Appendix F). Importantly, for this deterministic model, the calibrated errors ($\hat{q}$) are input-independent constants determined solely by the calibration data, representing global rather than instance-specific uncertainty.
@@ 3.6.2 Fine-tuned Model
We fine-tune the smaller MPP-AViT-Ti variant on MHD density fields using 75% of the data for training/calibration and 25% for validation. Here, the training data serves a dual purpose as calibration data, justified because fine-tuning targets a specific distribution, and we only care about performance within that regime.
Figure 12: Fine-tuned MPP-AViT performance on MHD density. Rows show: ground truth, predictions, absolute error, and calibrated error bar width (95% coverage). Fine-tuning dramatically improves accuracy compared to zero-shot (Figure 11). For this well-fitting deterministic model, calibrated errors capture global uncertainties in regions of high dynamics.
Figure 12 demonstrates substantial improvement over zero-shot performance. The fine-tuned model accurately captures both major and fine-scale density features. CP provides guaranteed coverage (Appendix F), though the error bars reflect global rather than local uncertainties due to the model’s deterministic nature. For well-fitted deterministic models like this, CP identifies regions of generally high dynamics across the dataset rather than instance-specific failure modes. Probabilistic models using the STD method would provide input-dependent bounds that adapt to each specific prediction.
@@ 3.7 Neural Weather Prediction
In addition to surrogate models of systems described by explicit PDEs, the proposed methodology is also applicable to more general machine learning models describing physical processes. To demonstrate this, we here study the use of CP for data-driven weather forecasting models. Traditional weather forecasting models typically combine PDEs describing large-scale interactions and parametrisations describing subgrid-scale physical processes (atmospheric_modeling_book). Data-driven machine learning models approximate this whole process with a single neural network model. This allows for orders of magnitude faster forecasting speed and, when training on data incorporating observations, also more accurate forecasts (ddw_risk; panguweather; pathak2022fourcastnet; lam2022graphcast).
Due to the chaotic nature of the weather system, capturing uncertainty in weather forecasts has long been an important consideration both in research and operations. Such probabilistic modelling has typically been achieved by ensemble forecasting, where perturbations are used to produce samples of possible forecast trajectories (fundamentals_of_nwp). Existing data-driven models are still largely deterministic (weatherbench2). There are attempts to produce ensemble forecasts using machine learning models by perturbing initial states (pathak2022fourcastnet; fuxi), training multiple models (calibration_of_large_neurwp) or generative modelling (swin_vrnn; price2024gencast; neural_lam). Fundamentally, ensemble forecasting always requires a computational cost proportional to the number of ensemble members, i.e. the number of forecasts made via perturbations. In contrast, CP offers a cheap method to immediately quantify forecast uncertainty for a time, position, and variable of interest. This uncertainty can be used by meteorologists interpreting the forecast, conveyed to decision-makers reacting to extreme weather events or directly presented to end-users looking up the forecast for the coming week. As CP enables UQ for a single forecast output by the model, it is directly applicable to existing deterministic machine learning models. A limitation of scalar uncertainty estimates is that there are no samples of the distribution over the atmospheric state. In some scenarios, it can be valuable to inspect such samples to gain an understanding of how different weather scenarios are unfolding.